A mixture of MnSO4 and MnSO4 x 4H2O has a mass of 2.005 g. After heating to drive off all the water the mass is 1.780 g. How many grams of MnSO4x4H2O are presented in the mixture? what is the mass prevent of MnSO4 x 4H2O in the mixture?
Answer #1

Mass of mixture=2.005 g

Mass remaining after driving off all the water=1.780 g

That means that the mass of MnSO4 remaining after driving off water=1.780 g

So mass of water=2.005 g - 1.780 g=0.225 g

Molar mass of water=2xMolar mass of H+Molar mass of O

=2x1 g/mol+16 g/mol=2 g/mol+16 g/mol=18 g/mol

Number of moles of water driven off=Mass of water/molar mass

=0.225 g/18 g/mol=0.0125 mol

We can see from the formula MnSO4x4H2O that for every 4 mol H2O there is 1 mol MnSO4 associated with it

So for 1 mol H2O there will be (1/4) mol MnSO4 associated with it

For 0.0125 mol H2OO there will be (1/4)x0.0125 mol=0.003125 mol MnSO4 associated with it

Molar mass of MnSO4=Molar mass of Mn+Molar mass of S+4xMolar mass of O=54.9 g/mol+32 g/mol+4x16 g/mol=54.9 g/mol+32 g/mol+64 g/mol=150.9 g/mol

So mass of MnSO4.4H2O present in the mixture=number of moles of MnSO4 associated with waterx Molar mass of MnSO4+Mass of water driven off

=0.003125 mol x 150.9 g/mol+0.225 g

=0.472 g+0.225 g=0.697 g

Mass percent of MnSO4.4H2O in mixture

=(Mass of MnSO4.4H2O/Mass of mixture)x100

=(0.697 g/2.005 g)x100=34.76 %